Find the domain and range of the following exponential functions:


1. 𝑓(π‘₯) = (


1


2


)


π‘₯+1


6. 𝑓(π‘₯) = 2


π‘₯βˆ’1 + 1


2. 𝑓(π‘₯) = 3


π‘₯+2 7. 𝑓(π‘₯) = βˆ’5


π‘₯ βˆ’ 2


3. 𝑓(π‘₯) = βˆ’(5)


π‘₯ 8. 𝑓(π‘₯) = (


1


4


)


2π‘₯+1


+ 3


4. (π‘₯) = 4


βˆ’π‘₯ 9. 𝑓(π‘₯) = 5


3π‘₯ βˆ’ 4


5. (π‘₯) = (


1


3


)


βˆ’π‘₯+1


10. 𝑓(π‘₯) = βˆ’ (


2


3


)


π‘₯


βˆ’ 3


1
Expert's answer
2021-11-12T14:48:32-0500

1. "f(x)=(\dfrac{1}{2})^x"

Domain: "(-\infin, \infin)"

Range: "(0, \infin)"


2. "f(x)=3^{x+2}"

Domain: "(-\infin, \infin)"

Range: "(0, \infin)"


3. "f(x)=-(5)^{x}"

Domain: "(-\infin, \infin)"

Range: "(-\infin, 0)"


4. "f(x)=4^{-x}"

Domain: "(-\infin, \infin)"

Range: "(0, \infin)"


5. "f(x)=(\dfrac{1}{3})^{-x+1}"

Domain: "(-\infin, \infin)"

Range: "(0, \infin)"


6. "f(x)=2^{x-1}+1"

Domain: "(-\infin, \infin)"

Range: "(1, \infin)"


7. "f(x)=-5^{x}-2"

Domain: "(-\infin, \infin)"

Range: "(-\infin, -2)"


8. "f(x)=(\dfrac{1}{4})^{2x+1}+3"

Domain: "(-\infin, \infin)"

Range: "(3, \infin)"


9. "f(x)=5^{3x}-4"

Domain: "(-\infin, \infin)"

Range: "( -4, \infin)"


10. "f(x)=-(\dfrac{2}{3})^{x}-3"

Domain: "(-\infin, \infin)"

Range: "(-\infin, -3)"